# The underlying structures of self-attention: symmetry, directionality, and emergent dynamics in Transformer training

### v1

@@ Acknowledgments
Let $U=[t_{0},\dots,t_{N}]$ a sequence of tokens, and let $t^{*}$ be the context of every token $t_{i}\in U$. Let $\{\mathbf{x}_{i}\}$ be the set of token embeddings associated with $U$ and let $\mathbf{y}$ be the token embedding of $t^{*}$. It follows from Proposition S1.7 that the squared norm of the weight update of the $k$-th column is given by
$\Delta\mathbf{w}_{\cdot,k}=\sum_{j\in\{t^{*}\}}\sum_{i\in P_{j}}\beta_{ij}\mathbf{x}_{i}[\mathbf{x_{j}}]_{k}=\sum_{i=1}^{N}\beta_{it^{*}}\mathbf{x}_{i}[\mathbf{y}]_{k}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{\cdot,k}||^{2}=[\mathbf{y}]^{2}_{k}||\sum_{i=1}^{N}\beta_{it^{*}}\mathbf{x}_{i}||^{2}\,,$ (S54)
while the squared norm of the weight update of the $m$-th row by
$\Delta\mathbf{w}_{m,\cdot}=\sum_{j\in\{t^{*}\}}\sum_{i\in P_{j}}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\mathbf{x}_{j}=\sum_{i=1}^{N}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\mathbf{y}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{m,\cdot}||^{2}=\big{(}\sum_{i=1}^{N}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\big{)}^{2}||\mathbf{y}||^{2}\,.$ (S55)
Let $\{\mathbf{x}_{i}\}\sim\mathcal{D}$ be a set of i.i.d. random vectors from a probability distribution $\mathcal{D}$ such that $\mathbb{E}[\mathbf{x}_{i}]=0$ and $\text{Cov}(\mathbf{x}_{i})=\Sigma\in\mathbb{R}^{d,d}$. Therefore, each $k$-th coordinate $[\mathbf{x}_{i}]_{k}$ is such that $\mathbb{E}[[\mathbf{x}_{i}]_{k}]=0$ and $\text{Var}([\mathbf{x}_{i}]_{k})=\Sigma_{k,k}$. We also assume that $\mathbf{y}$ is statistically independent from $\mathbf{x}_{i}\,\forall i$ with $\mathbb{E}[\mathbf{y}]=0$ and covariance $\text{Cov}(\mathbf{y})=\sigma_{y}^{2}\mathbb{I}$. It follows that the expected value of $||\Delta\mathbf{w}_{\cdot,k}||^{2}$ over $\mathcal{D}$ is given by
$\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]=\mathbb{E}\left[[\mathbf{y}]^{2}_{k}||\sum_{i=1}^{N}\beta_{it^{*}}\mathbf{x}_{i}||^{2}\right]=\mathbb{E}\left[[\mathbf{y}]^{2}_{k}\right]\mathbb{E}\left[||\sum_{i=1}^{N}\beta_{it^{*}}\mathbf{x}_{i}||^{2}\right]\,.$ (S56)
$\begin{split}\mathbb{E}\left[||\sum_{i=1}^{N}\beta_{it^{*}}\mathbf{x}_{i}||^{2}\right]&=\sum_{i,i^{\prime}}\mathbb{E}\left[\beta_{it^{*}}\beta_{i^{\prime}t^{*}}\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\sum_{i}\beta^{2}_{it^{*}}\mathbb{E}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i}\right]+\sum_{i^{\prime}\neq i}\beta_{it^{*}}\beta_{i^{\prime}t^{*}}\mathbb{E}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\sum_{i}\beta^{2}_{it^{*}}\text{Tr}\big{(}\mathbb{E}\left[\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right]\big{)}+\sum_{i^{\prime}\neq i}\beta_{it^{*}}\beta_{i^{\prime}t^{*}}\mathbb{E}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\text{Tr}(\Sigma)\sum_{i}\beta^{2}_{it^{*}}\,,\end{split}$ (S57)
and therefore
$\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]=\Gamma_{k,k}\text{Tr}(\Sigma)\sum_{i}\beta^{2}_{it^{*}}\,.$ (S58)
$\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]=\mathbb{E}\left[\big{(}\sum_{i=1}^{N}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\big{)}^{2}||\mathbf{y}||^{2}\right]=\mathbb{E}\left[\big{(}\sum_{i=1}^{N}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\big{)}^{2}\right]\mathbb{E}\left[||\mathbf{y}||^{2}\right]\,.$ (S59)
The first term can be decomposed as
$\mathbb{E}\left[\big{(}\sum_{i=1}^{n}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\big{)}^{2}\right]=\sum_{i=1}^{n}\beta^{2}_{it^{*}}\mathbb{E}\left[[\mathbf{x}_{i}]_{m}^{2}\right]+\sum_{i^{\prime}\neq i}\beta_{it^{*}}\beta_{i^{\prime}t^{*}}\mathbb{E}\left[[\mathbf{x}_{i^{\prime}}]_{m}[\mathbf{x}_{i}]_{m}\right]=\Sigma_{m,m}\sum_{i=1}^{n}\beta^{2}_{it^{*}}\,,$ (S60)
and therefore
$\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]=\Sigma_{m,m}\text{Tr}(\Gamma)\sum_{i}\beta^{2}_{it^{*}}\,.$ (S61)
The ratio of the expected value of these squared norms is
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{\Gamma_{k,k}\text{Tr}(\Sigma)}{\Sigma_{m,m}\text{Tr}(\Gamma)}=\frac{\Gamma_{k,k}}{\text{Tr}(\Gamma)}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}=\frac{1}{d}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}\,.$ (S62)
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{1}{d}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}>\frac{1}{d}\frac{d\,\Sigma_{m,m}}{\Sigma_{m,m}}=1\quad\forall m\,\,\text{s.t.}\,\,\Sigma_{m,m}<\frac{\text{Tr}(\Sigma)}{d}\,,$ (S63)
thus concluding the proof. ∎
Again, a complementary argument on asymmetric weight update can be made when a set of tokens $U=\{t_{1},\dots,t_{N}\}$ is used to predict a given token $t^{*}$, that is, $C_{i}=\{t^{*}\}\,\,\forall t_{i}\in U$. We formalize this in the following Corollary.
Corollary S1.9.
(Asymmetric growth of columns and rows for prediction). Let $U=[t_{0},\dots,t_{N}]$ a sequence of tokens, and let every token $t_{i}\in U$ be the context of a given token $t^{*}$. Let the associated embedding be drawn i.i.d. from a probability distribution $\mathcal{D}$ as in Proposition S1.8. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.2. Then the squared norm of the $m$-th row and $k$-th column of $\mathbf{W}_{qk}$ satisfies
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}<1\quad\forall m\in\{1,\dots,d\},\forall k\,\,\text{s.t.}\,\,\Sigma_{k,k}<\frac{\text{Tr}(\Sigma)}{d}$ (S64)
Proof.
Let $U=[t_{0},\dots,t_{N}]$ a sequence of tokens, and let $t^{*}$ be predicted by every token $t_{i}\in U$. Let $\{\mathbf{x}_{i}\}$ be the set of token embeddings associated with $U$ and let $\mathbf{y}$ be the token embedding of $t^{*}$. It follows from Proposition S1.7 that the squared norm of the weight update of the $k$-th column is given by
$\Delta\mathbf{w}_{\cdot,k}=\sum_{i=\{t^{*}\}}\sum_{j\in C_{i}}[\mathbf{x}_{j}]_{k}\mathbf{x}_{i}=\sum_{j=1}^{N}[\mathbf{x}_{j}]_{k}\mathbf{y}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{\cdot,k}||^{2}=\big{(}\sum_{i=1}^{N}[\mathbf{x}_{j}]_{k}\big{)}^{2}||\mathbf{y}||^{2}\,,$ (S65)
while the weight update of the $m$-th row follows
$\Delta\mathbf{w}_{m,\cdot}=\sum_{i=\{t^{*}\}}\sum_{j\in C_{i}}\mathbf{x}_{j}[\mathbf{x_{i}}]_{m}=\sum_{j=1}^{N}\mathbf{x}_{j}[\mathbf{y}]_{m}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{m,\cdot}||^{2}=[\mathbf{y}]^{2}_{m}||\sum_{i=1}^{N}\mathbf{x}_{j}||^{2}\,.$ (S66)
Therefore, following the same arguments as in Proposition S1.4.2, the ratio of the expected value of the square norms is given by
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{\Sigma_{k,k}\text{Tr}(\Gamma)}{\Gamma_{m,m}\text{Tr}(\Sigma)}=d\frac{\Sigma_{k,k}}{\text{Tr}(\Sigma)}<\frac{d\,\Sigma_{k,k}}{d\,\Sigma_{k,k}}=1\quad\forall k\,\,\text{s.t.}\,\,\Sigma_{k,k}<\frac{\text{Tr}(\Sigma)}{d}\,,$ (S67)
thus concluding the proof. ∎
S1.4.3 Average contribution of a given token for context and prediction
Additionally, we demonstrate that during autoregressive training, the expected number of tokens predicted by a specific token $t^{*}$ can differ from the expected number of tokens that predict $t^{*}$. More specifically, in autoregressive training, the ratio between these two expected numbers is influenced by the statistical correlations between tokens in the dataset. In contrast, for bidirectional training, the expected number of tokens predicted by $t^{*}$ and the number of tokens that predict $t^{*}$ are always equal. This equality holds true regardless of how tokens are correlated within the corpus, resulting in a ratio of 1. We formalize this in the following Proposition.
Proposition S1.10.
Let $\mathcal{V}=[t_{0},\dots,t_{V}]$ be a set of tokens. Let $\mathcal{U}$ be the sample space of all possible sequences of $N$ tokens, and let $U\in\mathcal{U}$ be a sequence $U=[t_{1},\dots,t_{N}]$. Let $\Pr[t_{j}=t^{*}]$ be the probability that the token at index $j$ in is given by $t^{*}\in\mathcal{V}$. Let $\mathbb{E}[\mu_{c}(t^{*})]$ be the expected number of tokens that are predicted by a given token $t^{*}$, and $\mathbb{E}[\mu_{p}(t^{*})]$ the expected number of tokens that predict a given token $t^{*}$. For autoregressive training, the ratio
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=\frac{\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]}{\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]}\,,$ (S68)
depends on $\Pr[t_{j}=t^{*}]$, while for bidirectional training the same ratio is given by
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=1\,.$ (S69)
Proof.
@@
Let $\mathcal{V}=[t_{0},\dots,t_{V}]$ be a set of tokens. Let $\mathcal{U}$ be the sample space of all possible sequences of $N$ tokens, and let $U$ be a sequence $U=[t_{1},\dots,t_{N}]\in\mathcal{U}$. Let $E=\{U\in\mathcal{U}:t_{k}=t^{*}\}$ the event of all sequences where the $k$-th token is $t^{*}$. The indicator function $\mathds{1}\{E\}(U)$ is a random variable defined as
$\mathds{1}\{E\}(U)=\begin{cases}1\,\,\text{if}\,\,t_{k}=t^{*}\\ 0\,\,\text{otherwise}\,,\end{cases}$ (S70)
and as such its expected value over $\mathcal{U}$ is
$\mathbb{E}\left[\mathds{1}\{t_{k}=t^{*}\}(U)\right]=\sum_{U\in\,\mathcal{U}}\mathds{1}\{t_{k}=t^{*}\}(U)\Pr[U]=\sum_{U\in\mathcal{U}\,:\,w_{j}=t^{*}}\Pr[U]=\Pr[t_{k}=t^{*}]\,.$ (S71)
@@
Let $\mu_{c}(t^{*})$ be the random variable quantifying the number of tokens predicted by $t^{*}$, while $\mu_{p}(t^{*})$ is the random variable quantifying the number of tokens predicted by $t^{*}$. We analyzed the case of autoregressive and bidirectional training separately.
@@
During autoregressive training, each time the token $t^{*}$ appears at position $k$ it is used as context to predict $N-k$ tokens, and it is predicted by $k-1$ tokens. It follows that $\mu_{c}(t^{*})$ is given by
$\mu_{c}(t^{*})=\sum_{l=2}^{N}\sum_{k=1}^{l-1}\mathds{1}\{t_{k}=t^{*}\}=\sum_{k=1}^{N}(N-k)\mathds{1}\{t_{k}=t^{*}\}\,,$ (S72)
while $\mu_{p}(t^{*})$ is given by
$\mu_{p}(t^{*})=\sum_{l=1}^{N-1}\sum_{k=1}^{l-1}\mathds{1}\{t_{l}=t^{*}\}=\sum_{k=1}^{N}(k-1)\mathds{1}\{t_{k}=t^{*}\}\,.$ (S73)
Therefore, the expected value of $\mu_{c}(t^{*})$ over all possible sequences is
$\mathbb{E}[\mu_{c}(t^{*})]=\mathbb{E}\left[\sum_{k=0}^{N}(N-k)\mathds{1}\{t_{k}=t^{*}\}\right]=\sum_{k=0}^{N}(N-k)\Pr[t_{k}=t^{*}]\,,$ (S74)
@@
$\mathbb{E}[\mu_{p}(t^{*})]=\mathbb{E}\left[\sum_{k=0}^{N}(k-1)\mathds{1}\{t_{k}=t^{*}\}\right]=\sum_{k=0}^{N}(k-1)\Pr[t_{k}=t^{*}]\,,$ (S75)
@@
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=\frac{\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]}{\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]}\,.$ (S76)
During bidirectional training, each time the token $t^{*}$ is masked at position $k$, it is predicted by $N-1$ tokens. We assume that the token $t^{*}$ significantly contributes to the context of a masked token only if it is itself not masked. During bidirectional training, whenever a token $t^{*}$ at position $k$ is masked, it is predicted using the other $N-1$ tokens in the sequence. We assume that $t^{*}$ significantly contributes to the context for predicting the masked token only if $t^{*}$ itself is not masked. Additionally, we assume that the probability $\rho$ of masking any given position $k$ is the same for all positions and does not depend on the specific token $t_{k}$ at that position. It follows that $\mu_{c}(t^{*})$ is given by
$\mu_{c}(t^{*})=\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{\text{mask}(k^{\prime})=1\}\,,$ (S77)
while $\mu_{p}(t^{*})$ is given by
$\mu_{p}(t^{*})=\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=1\}\sum_{k^{\prime}\neq k}\mathds{1}\{\text{mask}(k^{\prime})=0\}\,.$ (S78)
Therefore, their expected values are given by
$\begin{split}\mathbb{E}[\mu_{c}(t^{*})]&=\mathbb{E}\left[\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{\text{mask}(k^{\prime})=1\}\right]\\ &=\sum_{k=1}^{N}\mathbb{E}\left[\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{\text{mask}(k^{\prime})=1\}\right]\\ &=\sum_{k=1}^{N}\mathbb{E}\left[\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=0\}\right]\sum_{k^{\prime}\neq k}\left[\mathds{1}\{\text{mask}(k^{\prime})=1\}\right]\\ &=\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\text{Pr}(\text{mask}(k)=0)\sum_{k^{\prime}\neq k}\text{Pr}(\text{mask}(k^{\prime})=1)\\ &=N\rho(N-1)(1-\rho)\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\,,\end{split}$ (S79)
@@
$\begin{split}\mathbb{E}[\mu_{p}(t^{*})]&=\mathbb{E}\left[\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{\text{mask}(k)=1\}\sum_{k^{\prime}\neq k}\mathds{1}\{\text{mask}(k^{\prime})=0\}\right]\\ &=N\rho(N-1)(1-\rho)\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\,,\end{split}$ (S80)
@@
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=1\,,$ (S81)
thus concluding the proof.
Let us assume that there is no statistical correlation between the tokens in $U$, we can factorize the joint probability $\Pr[U]$ as follows
$\Pr[U]=\Pr[t_{1},\dots,t_{N}]=\Pi_{i=1}^{N}\Pr[t_{i}]\,.$ (S82)
Additionally, let us assume that each token is identically distributed independently of the position. Therefore, during autoregressive training, the expected number $\mathbb{E}[\mu_{c}(t^{*})]$ can be simplified as
$\mathbb{E}[\mu_{c}(t^{*})]=\sum_{k=0}^{N}(N-k)\Pr[t_{k}=t^{*}]=\sum_{k=0}^{N}(N-k)\Pr[t^{*}]=\Pr[t^{*}]\frac{N(N-1)}{2}\,,$ (S83)
and the same is true for $\mathbb{E}[\mu_{p}(t^{*})]$
$\mathbb{E}[\mu_{p}(t^{*})]=\sum_{k=0}^{N}(k-1)\Pr[t_{k}=t^{*}]=\sum_{k=0}^{N}(k-1)\Pr[t^{*}]=\Pr[t^{*}]\frac{N(N-1)}{2}\,.$ (S84)
It follows that, in the case of statistical independence between tokens, the expected number of tokens predicted by $t^{*}$ and the expected number of tokens predicting $t^{*}$. Under these assumptions, there is no difference between autoregressive and bidirectional training.
On the other hand, let us assume a general joint probability between tokens, $\Pr[U]=\Pr[t_{1},\dots,t_{N}]$. When conditioning the probabilities over the future tokens, we obtain a factorization of the joint probability as follows,
$\Pr[U]=\prod_{i=1}^{N}\Pr[t_{i}\mid t_{i+1},\dots,t_{N}]\,.$ (S85)
The probability of the $k$-th token being $t^{*}$ across the possible sequences is thus given by
$\begin{split}\Pr[t_{k}=t^{*}]&=\sum_{t_{1},\dots,t_{N}}\mathds{1}\{t_{k}=t^{*}\}\Pr[t_{1},\dots,t_{n}]\\ &=\sum_{\begin{subarray}{c}t_{1},\dots,t_{k-1},t_{k+1},\dots,t_{N}\end{subarray}}\left(\prod_{\begin{subarray}{c}l=1\\ l\neq k\end{subarray}}^{T}\Pr[t_{l}|t_{l+1},\dots,t_{N}]\right)\cdot\Pr[t_{k}=t|t_{k+1},\dots,t_{N}]\,,\end{split}$ (S86)
leading to the ratio
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=\frac{\sum_{k=1}^{N}(N-k)\sum_{\begin{subarray}{c}t_{1},\dots,t_{k-1},t_{k+1},\dots,t_{N}\end{subarray}}\left(\prod_{\begin{subarray}{c}l=1\\ l\neq k\end{subarray}}^{T}\Pr[t_{l}|t_{l+1},\dots,t_{N}]\right)\cdot\Pr[t_{k}=t|t_{k+1},\dots,t_{N}]}{\sum_{k=1}^{N}(k-1)\sum_{\begin{subarray}{c}t_{1},\dots,t_{k-1},t_{k+1},\dots,t_{N}\end{subarray}}\left(\prod_{\begin{subarray}{c}l=1\\ l\neq k\end{subarray}}^{T}\Pr[t_{l}|t_{l+1},\dots,t_{N}]\right)\cdot\Pr[t_{k}=t|t_{k+1},\dots,t_{N}]}\,.$ (S87)
Therefore, when $t^{*}$ has a high probability of occurring before a given set of tokens, or when $t^{*}$ is likely to occur at the beginning of sentences and specific pieces of text, this ratio is likely to be higher than 1.
S1.4.4 Theorem on the emergence of directionality
Here, we show that, under the same assumptions as in Proposition S1.8, the weight updates of $\mathbf{W}_{qk}$ induce column dominance during autoregressive training. Specifically, there are on average more columns with high norms than rows with high norms. We assume that, at initialization, the entries of $\mathbf{W}_{qk}$ are drawn from a probability distribution $\mathcal{P}$ with finite mean $\mu$ and variance $\sigma^{2}$. This assumption holds for any standard machine learning initialization scheme. This implies that each $k$-th column $\mathbf{w}_{\cdot,k}$ and $m$-th row $\mathbf{w}_{m,\cdot}$ have the same mean $\mu$ and variance $\sigma^{2}/\sqrt{n}$. Furthermore, the squared norm of both rows and columns have equal mean $n(\sigma^{2}+\mu^{2})$ and variance $n(2\sigma^{4}+4\mu^{2})$.

### v2

@@ Acknowledgments
Let $\mathcal{V}=\{t_{0},\dots,t_{V}\}$ be a set of tokens. Let $\mathcal{U}=\mathcal{V}^{N}$ denote the sample space of all possible sequences of $N$ tokens from the vocabulary $\mathcal{V}$. Let $\mathcal{F}=2^{\mathcal{U}}$ be the $\sigma$-algebra given by the power set of $\mathcal{U}$, and let $P$ be the probability distribution over $\mathcal{U}$ that defines the distribution of the random variable sequence $U$. This defines the probability space $(\mathcal{U},\mathcal{F},P)$. For each position $k\in\{1,\dots,N\}$ we let the random variable $T_{k}$ be $T_{k}:\mathcal{U}\rightarrow\mathcal{V}$ such that $T_{k}(U)\equiv t_{k}$. Let $E=\{U\in\mathcal{U}:t_{k}=t^{*}\}$ the event of all sequences where the $k$-th token is a fixed token $t^{*}$. The indicator function $\mathds{1}\{E\}(U)$ is a random variable defined as
$\mathds{1}\{E\}(U)=\mathds{1}\{t_{k}=t^{*}\}(U)=\begin{cases}1\,\,\text{if}\,\,t_{k}=t^{*}\\ 0\,\,\text{otherwise}\,,\end{cases}$ (S64)
and as such its expected value over $P$ is
$\mathbb{E}_{P}\left[\mathds{1}\{t_{k}=t^{*}\}(U)\right]=\sum_{U\in\,\mathcal{U}}\mathds{1}\{t_{k}=t^{*}\}(U)\Pr[U]=\sum_{U\in\mathcal{U}\,:\,t_{k}=t^{*}}\Pr[U]=\Pr[t_{k}=t^{*}]\,.$ (S65)
@@
Let $\mu_{c}(t^{*},U)$ be the random variable quantifying the number of tokens predicted by $t^{*}$, while $\mu_{p}(t^{*},U)$ is the random variable quantifying the number of tokens predicted by $t^{*}$. We analyzed the case of autoregressive and bidirectional training separately.
@@
During autoregressive training, each time the token $t^{*}$ appears at position $k$ it is used as context to predict $N-k$ tokens, and it is predicted by $k-1$ tokens. It follows that $\mu_{c}(t^{*},U)$ is given by
$\mu_{c}(t^{*},U)=\sum_{l=2}^{N}\sum_{k=1}^{l-1}\mathds{1}\{t_{k}=t^{*}\}=\sum_{k=1}^{N}(N-k)\mathds{1}\{t_{k}=t^{*}\}\,,$ (S66)
while $\mu_{p}(t^{*},U)$ is given by
$\mu_{p}(t^{*},U)=\sum_{l=1}^{N-1}\sum_{k=1}^{l-1}\mathds{1}\{t_{l}=t^{*}\}=\sum_{k=1}^{N}(k-1)\mathds{1}\{t_{k}=t^{*}\}\,.$ (S67)
Therefore, the expected value of $\mu_{c}(t^{*},U)$ over $P$ is
$\mathbb{E}_{P}[\mu_{c}(t^{*},U)]=\mathbb{E}_{P}\left[\sum_{k=1}^{N}(N-k)\mathds{1}\{t_{k}=t^{*}\}\right]=\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]\,,$ (S68)
@@
$\mathbb{E}_{P}[\mu_{p}(t^{*},U)]=\mathbb{E}_{P}\left[\sum_{k=1}^{N}(k-1)\mathds{1}\{t_{k}=t^{*}\}\right]=\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]\,,$ (S69)
@@
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=\frac{\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]}{\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]}\,.$ (S70)
During bidirectional training, each time the token $t^{*}$ is masked at position $k$, it is predicted by $N-1$ tokens. We assume that the token $t^{*}$ significantly contributes to the context of a masked token only if it is itself not masked (see Appendix S1.3). Additionally, we assume that the probability $\rho$ of masking any given position $k$ is the same for all positions and does not depend on the specific token $t_{k}$ at that position. As such, we define masking as an independent Bernoulli variable $m_{k}:\Sigma_{k}\rightarrow\{0.1\}\sim P_{m}$ such that $\text{Pr}[m_{k}=1]=\rho$, independently of the token and sequence spaces. It follows that $\mu_{c}(t^{*},U,m)$ is given by
$\mu_{c}(t^{*},U,m)=\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{m_{k^{\prime}}=1\}\,,$ (S71)
while $\mu_{p}(t^{*},U,m)$ is given by
$\mu_{p}(t^{*},U,m)=\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=1\}\sum_{k^{\prime}\neq k}\mathds{1}\{m_{k^{\prime}}=0\}\,.$ (S72)
Therefore, by taking the expectation with respect to both distributions $P$ and $P_{m}$, their respective expected values are given by
$\begin{split}\mathbb{E}_{P,P_{m}}[\mu_{c}(t^{*},U,m)]&=\mathbb{E}_{P,P_{m}}\left[\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{m_{k^{\prime}}=1\}\right]\\ &=\sum_{k=1}^{N}\mathbb{E}_{P,P_{m}}\left[\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=0\}\sum_{k^{\prime}\neq k}\mathds{1}\{m_{k^{\prime}}=1\}\right]\\ &=\sum_{k=1}^{N}\mathbb{E}_{P,P_{m}}\left[\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=0\}\right]\sum_{k^{\prime}\neq k}\left[\mathds{1}\{m_{k^{\prime}}=1\}\right]\\ &=\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\text{Pr}(m_{k}=0)\sum_{k^{\prime}\neq k}\text{Pr}(m_{k^{\prime}}=1)\\ &=N\rho(N-1)(1-\rho)\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\,,\end{split}$ (S73)
@@
$\begin{split}\mathbb{E}_{P,P_{m}}[\mu_{p}(t^{*},U,m)]&=\mathbb{E}_{P,P_{m}}\left[\sum_{k=1}^{N}\mathds{1}\{t_{k}=t^{*}\}\mathds{1}\{m_{k}=1\}\sum_{k^{\prime}\neq k}\mathds{1}\{m_{k^{\prime}}=0\}\right]\\ &=N\rho(N-1)(1-\rho)\sum_{k=1}^{N}\text{Pr}(t_{k}=t^{*})\,,\end{split}$ (S74)
@@
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=1\,,$ (S75)
where we omit the dependence on the independent random variable $m$ for simplicity, thus concluding the proof. ∎
The following remark highlights how, during autoregressive training, the ratio of expected token counts can vary depending on the correlations and conditional dependencies among the tokens.
Remark S1.10.
We separately analyze the following two cases:
  • No statistical correlation between tokens. Assume that the tokens are statistically independent and identically distributed (i.i.d.) across positions, i.e., the probability of each token is the same at every position and does not depend on surrounding tokens. Under this assumption, during autoregressive training, the expected number of times token $t^{*}$ is used to predict future tokens is given by,
    $\mathbb{E}_{P}[\mu_{c}(t^{*},U)]=\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]=\Pr[t^{*}]\sum_{k=1}^{N}(N-k)=\Pr[t^{*}]\frac{N(N-1)}{2}\,.$ (S76)
    Similarly, the expected number of times token $t^{*}$ is predicted by preceding tokens is,
    $\mathbb{E}_{P}[\mu_{p}(t^{*},U)]=\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]=\Pr[t^{*}]\sum_{k=1}^{N}(k-1)=\Pr[t^{*}]\frac{N(N-1)}{2}\,.$ (S77)
    Therefore, when tokens are independent and identically distributed, the expected number of tokens predicted by $t^{*}$ and the expected number of tokens used to predict $t^{*}$ are equal. In this case, there is no difference between autoregressive and bidirectional training.
  • Statistical correlation between tokens. Let us assume that the probability distribution $P$ encodes statistical dependencies between tokens. Define $\mu(t^{*})$ as the expected position index of token $t^{*}$,
    $\mu(t^{*})=\frac{\sum_{k=1}^{N}k\Pr[t_{k}=t^{*}]}{\sum_{k=1}^{N}\Pr[t_{k}=t^{*}]}\,.$ (S78)
    Using this, the ratio of expected counts under autoregressive training becomes,
    $\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=\frac{N-\mu(t^{*})}{\mu(t^{*})-1}\,.$ (S79)
    This expression shows that if $t^{*}$ tends to appear earlier in the sequence, that is, if $\mu(t^{*})<(N+1)/2$ (left of center), then the ratio is greater than 1. Conversely, if $t^{*}$ appears later in the sequence (right of center), the ratio is less than 1. More formally, suppose there exists $\delta>0$ such that
    $\mu(t^{*})\leq\frac{N+1}{2}-\delta\,.$ (S80)
    Then the ratio can be rewritten as,
    $\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*}),U]}-1=\frac{N+1-2\mu(t^{*})}{\mu(t^{*})-1}\,,$ (S81)
    where the numerator satisfies $N+1-2\mu(t^{*})\geq 2\delta$, and the denominator is positive for any token that can appear at position 1 (i.e., $\mu(t^{*})>1$). Hence, the ratio is bounded below by,
    $\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}\geq 1+\varepsilon(t^{*})\,,\quad\text{with}\quad\varepsilon(t^{*})=\frac{2\delta}{\mu(t^{*})-1}\,.$ (S82)
S1.5.4 Tail probabilities of distributions with equal mean and different variances